OpenStudy (anonymous):

Solve: 4(x + 1) = −12

OpenStudy (anonymous):

4(x+1)=-12 4x+4=-12 4x=-16 x=-4

OpenStudy (anonymous):

4x+4=-12 4x = -12-4 4x = -16 x=-16/4 x=-4

OpenStudy (anonymous):

4x+4 = -12 (expanded equation) 4x+4-4= -12-4 (subtract 4 from both sides) 4x=-16 4x/4 = -16/4 (divide both sides by 4) x = -4